OpenStudy (anonymous):

Find the standard equation and graph of a parabola that matches the given set of characteristics Vertex (0,0) Focus (4,0)

OpenStudy (sidsiddhartha):

standard equation can be written as (y-k)^2=4p(x-h) where h,k is the coordinate of the vertex

OpenStudy (sidsiddhartha):

now distence between focus and vertex is=16 so 4p=16 p=4 so equation of the parabola y^2=4px Y^2=16x it will look like